solve the equation: in e^x-In e^4=In e^5
Answer by Alan3354(69443)

solve the equation: in e^x-In e^4=In e^5
--------------
I think you mean ln, natural log
ln e^x-In e^4=In e^5
x - 4 = 5
x = 9